### v2

# Low Resource Reconstruction Attacks Through Benign Prompts

Abstract
The recent advances in generative models such as diffusion models have raised several risks and concerns related to privacy, copyright infringements and data stewardship. To better understand and control the risks, various researchers have created techniques, experiments and attacks that reconstruct images, or part of images, from the training set. While these techniques already establish that data from the training set can be reconstructed, they often rely on high-resources, excess to the training set as well as well-engineered and designed prompts.
In this work, we devise a new attack that requires low resources, assumes little to no access to the actual training set, and identifies, seemingly, benign prompts that lead to potentially-risky image reconstruction. This highlights the risk that images might even be reconstructed by an uninformed user and unintentionally. For example, we identified that, with regard to one existing model, the prompt “blue Unisex T-Shirt” can generate the face of a real-life human model. Our method builds on an intuition from previous works which leverages domain knowledge and identifies a fundamental vulnerability that stems from the use of scraped data from e-commerce platforms, where templated layouts and images are tied to pattern-like prompts. [^1] Our attack implementation is publicly available at https://github.com/TheSolY/lr-tmi.
1 Introduction
With the increasing use of generative and foundation models, and the increasing use of public and potentially private data for training, a range of new risks and challenges has evolved within the field of Machine Learning. Today, there is growing concern that the use of data may lead to breaches of copyright and violations of individual privacy. While privacy is typically associated with sensitive or non-public data, even public data raises important concerns [34]. Individuals who have consented to share their photos publicly may still expect control over how their data is used, particularly to prevent it from being presented in unintended or inappropriate contexts.
Source Generated Prompt
Refer to caption Refer to caption "Galaxy Area Rug"
(source: GearFrost)
Refer to caption Refer to caption "Floral iPhone Case"
(source: eBay)
Refer to caption Refer to caption "Abstract Art Shower Curtain"
(source: Amazon)
Figure 1: Examples of template memorized images reconstructed through our attack on SD 1.4.
In turn, researchers are now investigating how data is memorized within generative and foundation models, and whether data can be extracted in ways unintended. Several works [33, 13, 36] demonstrated that attacks can be designed to extract, blatantly and verbatim, data that appeared in the training sets. These existing attacks tend to rely on access to the training data, some on substantial computational resources, and they search for specific prompts, from training data that, potentially, trigger the extraction. In that sense, these attacks simulate a malicious adversary that explicitly aims to extract such data. However, an important concern is the potential for unintentional image conjuring, where a user issues a benign prompt that might inadvertently trigger the same phenomenon.
Thus, towards a better understanding of the potential risks, we develop in this work a followup attack. Our objective is to construct simple and benign prompts that call for generic objects such as t-shirt, round metal wall art, beach towel etc., without too-specific details. We applied our attack on previously attacked model, and use these prompts to extract images that present elements and objects whose source can be traced to other websites. Perhaps the most disturbing feature of our attack is that real-life models, whose images appeared in such websites, are also extracted by these, so-called innocent prompts, as seen in Fig. 4. In comparison, previous attacks could extract real images of real people but typically relied on intentional prompts, and the scope was to generate the exact image. For instance, Carlini et al. [13] showed that prompting with “Ann Graham Lotz” could yield a verbatim copy of an existing image. The objective was not to show that her image will be extracted (as the prompt requests that) but to investigate the copying of the existing photo. In contrast, our attack shows that even unintentional prompts can generate images of real individuals without explicitly requesting them. This behavior raises distinct concerns around privacy, and individuals’ rights. Particularly the right not to have their likeness used for modeling purposes without their consent.
Previous work focused on understanding the training data, and in identifying duplications. Our attack avoids harvesting the training data and builds on a working hypothesis that data from e-commerce sites and Print on Demand websites (PoDs) is (intentionally or unintentionally) harvested during training. Then, we design an attack that leverages specific traits and domain-knowledge regarding e-commerce sites.
PoD platforms use automated design placement tools that overlay artwork onto pre-existing images of product by using smart masks and blending techniques. A single platform output may be integrated into many other websites, and the system can instantly generate realistic previews showing the design on different product types, angles, and lighting conditions. As a result, such websites display many images that are identical up to a fixed region where the design is placed. The LAION-5B dataset [31], a large-scale, web-scraped image-text dataset, likely contains a significant number of PoD-generated images due to their widespread presence on the internet. We could validate this claim through existing datasets of duplicated images on LAION, [37]. Importantly though, because generated images are not verbatim copies, we could also validate that many variants of the same image were not necessarily flagged as duplications by these identification systems. In other words, these platforms produce images that may consistently feature repeated elements, such as the same design or product, but with variations in context, angle, background, or lighting. In turn, the images are not always deemed as duplicates, even though they share substantial visual content.
Discussion and Limitations
Extracting training data presents a major challenge, raising serious concerns around privacy and potential copyright infringement. Our results focus on the phenomenon of unintentional extraction of memorized data. Our attack demonstrates that image memorization in diffusion models can be exploited with minimal resources and without access to the training dataset, posing a more widespread privacy and copyright risk than previously understood. Because our attack relies on seemingly innocuous prompts (e.g., "Abstract Art Unisex T-Shirt") we reveal that users may unintentionally generate memorized images of real individuals or copyrighted content. These types of attacks amplify the risks associated with inadvertent copying and highlight the troubling possibility that even seemingly benign users may inadvertently trigger such leaks.
The vulnerability we exploit arises from the template-structure of scraped e-commerce data, where uniform layouts across product categories facilitate memorization and make leakage more detectable. While engineers already attempt to clean datasets of obvious duplicates, our attack suggests that duplication should be interpreted in a more conservative sense, that also accounts for repeated patterns and templates. Most of our attacks were conducted on the SD 1.4 version, but even in the most recent models such as Stable Diffusion v3.5, Flux Schnell, and Midjourney v6, we could still obtain some of the results from version 1.4, as presented in Section 4.2. We could also trace some memorization of previously reconstructed images, as presented in Appendix B. It is possible that methods of decoupling between text and image that were presented in the training process of version 3.5 significantly strengthened the model against the threat of our attack, but it did not completely mitigate it. The examples for successful extraction of template-memorized images from version 3.5 which are shown in Fig. 6 hint that it is not completely immune to our attack. We leave it to future study to further improve our attack, as well as to develop more principled methods to protect models from similar attacks.
As we discuss, the ability to extract memorized data poses various potential risks. In this work we focused solely on extracting data, or variants thereof, that is already accessible. Specifically, all images of identifiable individuals that are presented in this work are limited to cases where the original image is publicly available, easily accessible on the internet and the alterations are limited to benign visual modifications—such as changes in clothing color or background.
Related Work
In recent years several reconstruction attacks for foundation models have been designed, including attacks targeted on language models [12, 25, 24] and, more aligned with out work, image reconstruction [33, 13, 36]. We elaborate further on these attacks and how they relate to our work in Section 2.
We are specifically concerned with attacks that generate data from prompts, or black-box (to various extent) models that through the “standard” use of the model induce types of image conjuring. This is distinct from various attacks that reconstruct training data by analyzing and probing the model [19, 39, 17]. These works highlight the memorization of training data by large learning models. Such memorization is necessary for reconstruction attacks to even be possible. There is increasing theoretical support that such memorization may be unavoidable [26, 10, 35, 16], and reconstruction attacks complement these works by showing that such memorization has implications. This leads to the open problem of mitigating risks and prevent data reconstruction even when the model inherently retains information from its training data.
Given that data is memorized, an important aspect that arises is the question of originality, or how to make sure that the output of the model is original and not mere interpolation of existing data. This leads to an interesting theoretical question – to formalize and model what constitute original data [15, 30], and to an imminent practical question – to identify and regulate non-originality [20, 18, 14]. Our attack generates results in a middle ground between blatant, verbatim copying and non-copying that lacks originality, where the system replicates elements without directly copying. Such types of interpolations have been observed before [9, 32]. It is possible that such interpolations can be mitigated and regulated through appropriate credit attribution. The question of how to properly attribute credit, though, is an ongoing research question [27].
Our attack also highlight the risks that can be further amplified through jail-breaking methods [28, 38, 40]. Jail breaking attacks don’t necessarily try to reconstruct data but instead to generate inappropriate usage. We did not attempt to conduct such experiments. However, our attacks generated existing objects and people in new contexts and backgrounds, taken together with such jail-breaking attacks the potential risk should be considered.
2 Background and Existing Attacks
The focus of this work is on a new low-resource attack that avoids access to the training set, harvesting data duplications or have knowledge of prompts used in training. While we devise a methodology that can circumvent the use of such access, the intuition behind our attack builds on insights and techniques that were gathered in previous work. Therefore, we briefly review previous works and attacks as well as the important intuition that we build from them.
Random selection
Somepalli et al. [32] demonstrated that by analyzing a random subsample of trained data, replicated content can be generated quite straightforwardly. After designing a detection system for replicated content, they randomly sample 9000 source images from a $12M$ subset of images that were also used in the final rounds of training. Their detection system identified replicated content using the prompts from the training dataset. Interestingly, captions from the dataset don’t typically generate the training image, but may conjure a memorized image. This attack relies on a small subset on which the model was further trained, and where a typical random sample is duplicated roughly $11$ times [32]. Then the attack generates specific captions that appeared in the dataset verbatim, but they hypothesize that certain key phrases can indeed cause unintentional replicated content. Our work validates this claim and indeed builds on less detailed prompts and short key phrases whose origin were not extracted from the training set, making them much more likely to be used unintentionally.
Training data duplication
OpenAI [29] collected a dataset of prompts that often resulted in duplicated generated images on their model. They generated images from them, and found the closest in perceptual similarity to an image in the training set. They demonstrated how memorized images had many near-duplicates in the training set.
Carlini et al. [13] reasons that training data duplication is a potential cause for memorization and presents the hypothesis that images extracted from memorization, as opposed to novel generated data, will also contain near-duplicates. Their attack takes the most frequently duplicated images along with their captions, generates many images from each caption, and then looks for cliques of nearly identical images. Cliques of at least 10 images are predicted as memorized. The clique search uses $\ell_{2}$ similarity on patches, similar to the split-product method in [32], and the training data duplicate search relies on entire-image $\ell_{2}$ and CLIP embedding. Overall, this attack is resource-intensive and depends on broad access to the training dataset. In addition, the attack is focused on Verbatim copying. Searching for replicated content, we often look for objects in the background or foreground [32]. Thus, an image may be generated only once and still contain replicated content. Identifying such content by inspecting cliques, though, can be challenging.
One-Step Synthesis
Webster [36] takes a different theoretical approach, selecting the candidates for memorization based on their one-step synthesis behavior, under the hypothesis that memorized image-text pairs present clear, consistent edges after the first denoising step, while non-memorized pairs are blurry after the first denoising step. Using this method they indeed found many captions that extract template memorized images.
To select the candidate captions they relied on full duplicate of image-text pairs. The candidates were selected as the highest scoring 2 million image-text pairs from LAION2B on the duplication metric presented in [37]. The 2 million candidate include only fully duplicated text-image pairs, and does not include images that were duplicated in a partial fixed region, along with a partial text, but for which no full duplication existed. Out of the 2M candidates, 30k were selected based on their score in the white-box attack presented at [36], i.e. those for which the most noise had been removed at the first denoising step of Stable Diffusion V1. Due to the method of identifying duplicates, they identified a phenomena of template memorization, where only a certain object in the image is being duplicated. They are also the first to postulate that such images might be traced to e-commerce sites, a fact that we further validate and helped us in devising our attack that does not require to harvest duplications from the training dataset. Therefore, we expand on the phenomena of template memorization, showing that one could use a natural English text including a short Collocation rather than a highly specific text, and that such Collocations can extract several image templates rather than a specific one. We further rely on e-commerce websites to select candidate prompts and collocations, releasing us from searching and using full duplicates, thus expanding the search to image-text pairs that were duplicated only in part.
3 Our Attack
3.1 Data Collection
Our attack, as in previous work, first builds on extracting potential prompts that will trigger copied works. A distinctive trait of our attack is that we avoid any access to the training set, and want to generate more natural prompts that will lead to reconstructed data. Because we avoid access to the training set, we cannot perform the partial-duplication search described. Instead, we collect a list of candidate expressions which are associated with e-commerce websites that were scraped from LAION-5B. We started with an arbitrary list of such sites by asking ChatGPT through the prompt “consumer facing print on demand websites that appear in LAION 5B.”. Then, we scraped a list of categories and potential sold objects. Other ways to choose the websites ended with similar lists.
For each of these sites, we look at the snapshot of the site from March 2021, since the latest Common Crawl version to enter LAION 5B was April 2021. We then scraped a list of collocations from categories that are sold on these websites such as “unisex t-shirts", “athletic shoes" etc. Overall we reached a list of around $108$ collocations. In e-commerce websites, images for each category are generated through different templates. Therefore, in order to generate diversity, that will allow us to identify duplicates, we generated, for each collocation, a set of prompts by adding a description-visual pattern to each. The descriptions were: "Galaxy", "Floral, "Abstract Art" , "I Heart ML", "blue", "red". This process leads to a set of prompts, pattern+collocation phrases, such as “Floral Unisex t-shirts". For each such prompt we generate $25-50$ different images (distinct by their seeds from 00 to $49$).
Collections from previous work
Only for the sake of comparison with previous work, we also extracted list of categories that were related to sources that previous work also identified. These categories allow us to further analyze our attack and enable comparisons. In more detail, we took the prompts already identified by Webster [36] as what they denote "templates verbatim". These are images memorized up to a fixed mask, or "retrieval verbatim"222these are identified by generated images in a near duplicate of a training image but the caption of the training image is different than the prompt used to generate the generated image.. Similarly we extracted collocations from Somepalli et al. [32], Hintersdorf et al. [21] that we identified as product descriptions (rather than news titles etc.). From these, we manually isolated prompted collocations as the 1-3 words long phrase describing a product. For example, in the list of prompts memorized by Stable Diffusion, published by Webster [36], appears the prompt "Outshak Hand-Knotted Beige/Blue Area Rug by Shalom Brothers" out of which we extracted the collocation "Area Rug". After collecting a list of candidate expressions, we generated many images similar to our previous method where we add to the collocation a description pattern.
3.2 Searching near-duplicates
Similarly to previous works [29] [13], we search for near duplicates among the image generated through the process depicted above. One advantage of our method is that generated images are from known categories. This allows us to use image segmentation pre-trained models to predict the mask of the editable region and filter it out prior to the near-duplicates search. For household items we used [8], and for clothing we used [6].
We then run clique search based on the similarity of the predicted fixed region, where adjacent images are pairs with cosine similarity of above 0.95 on the CLIP embedding. Clique of at least size 2 were suspected as template-memorized, and were filtered to identify near-duplicates. In addition, using CLIP instead of $\ell_{2}$ distance, included images with perturbations which are minor visual variations between images (see Section 5.2), that are otherwise very similar in the same clique. We also searched for near-duplicates through manual visual inspection to complete the results missing from the limitation of the segmentation models.
“X Area Rug” (“rug”) Refer to caption
"X Wall Tapestry" ("painting") Refer to caption
“X Shower Curtain” (“curtain”) Refer to caption
“X High-Tops Sneakers” (“Right-shoe”) Refer to caption
Figure 2: Examples of image cliques found through our segmentation and masking method. Segmentation category is listed in brackets.
3.3 Tracing the Source Image
While for duplicated images we suspect that they are generally copied, for several of the images, we could validate and find a source. Sources were traced by Google Lens search, by targeting search to e-commerce websites and through visual inspection of images associated with the website categories.
As we could narrow down the search of the source to handful number of websites, some images were recognized even without duplications, such as the generated image shown in first row of Fig. 4. For this reason, we then continued to visually inspect the generated images comparing them to the original images in the origin websites, under their product categories, leading to exposing of more such memorized images.
4 Results
All of our experiments were conducted on a single RTX A6000 Machine. We primarily focused on Stable Diffusion version 1.4 from the official checkpoint accessible at HuggingFace [4], which is the model where previous attacks were introduced [13, 36, 33]. Results from more recent models are discussed and reported in Section 4.2.
In Figure 2 we demonstrate examples of four cliques we identified in four different collocations. Since, we generate the cliques using CLIP embedding and not $\ell_{2}$ distance, notice that in the curtain example cliques may include slightly zoomed-in versions of the same image. On the other hand, some of the duplicates (e.g. the rugs) can differ only in the color of the rug (that is being segmented and masked for clique identification), while others may involve differing patterns (for example, the wall tapestries).
4.1 Images with identified source
For many of the duplicated images, we could locate a source image from the web. Some examples are depicted in Fig. 3. Noteworthy, even for some images that we did not identify them as duplicated, we could identify a source image by inspecting e-commerce sites from which we’ve taken the categories, for example the image on the first row of Fig. 4.
Category Abstract Art Galaxy Floral I Heart ML Source
Beach Towel Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Throw Pillow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T-Shirt Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T-Shirt Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Examples across product categories.
In Fig. 8, we show reconstructed images matched LAION sources and their captions, even though the prompts used were generic and the captions themselves. This highlights the risk of unintentionally reproducing training data.
Real Humans
Arguably, the most concerning type of image recall are those which potentially contain a real human model. Several of the duplicates we identified indeed contain real humans that we could find their source, see Fig. 4.
As discussed, previous works also managed to extract real-humans [13]. However this was generally done by prompts that actively request that person’s image, and the focus was on copyright and verbatim copy. Here, the person is extracted by a prompt that, seemingly, does not contain their name, but instead the prompt contains a collocation associated with a product that the model had presented in the image. For example, mundane items such as "T-Shirt". Even in cases where the face was deformed (as often happens in SD) or the picture is only of the torso, it might still contain identifying markers such as tattoos as well as other specific visual attributes such as haircut and pose. This is exemplified in Fig. 9 (left columns).
Source & Generated Image Prompt Used & Source
Refer to caption “blue Unisex T-Shirt”
(source: RedBubble)
Refer to caption “Abstract Art A-Line Dress”
(source: RedBubble)
Refer to caption “Abstract Art T-Shirt”
(source: Zazzle)
Figure 4: Examples of generated people whose source images were identified (left: source, right: generated). The first row shows a person identified by name [11].
4.2 Attacks on State-Of-The-Art and other models
While we focused out efforts, and designed our attack on Stable Diffusion 1.4 (SD 1.4), we could also extract template-memorized images from other popular generative models. Primarily, by using the same collocations identified in SD 1.4. In Fig. 5 we exhibit further images extracted from DeepFloyd IF-XL-I-v1.0 [1] and Midjourney V4 [3]. Both models were released around the time of SD 1.4, and in Fig. 6 we demonstrate further extractions from state-of-the-art (SOTA) models.
  • DeepFloyd IF-XL-I-v1.0: although the generated images exhibited lower contrast, which made clique search impractical, we successfully identified template-memorized images through visual inspection. In addition, we discovered new templates—absent from SD 1.4—using Google Lens searches on manually flagged generations.
  • Midjourney V4: MidJourney [3] posed additional challenge for our attack, due to the lack of an open-source implementation or public API. However, by restricting our experiments to a limited set of prompts and seeds, selected based on template-memorized prompts identified in SD 1.4, we could still apply our attack.
    For each prompt and seed, we collected the four default outputs returned by Midjourney, specifying version 4 with the argument --v 4. Despite limited prompt and seed control, we successfully extracted template-like generations from Midjourney. Our results included both direct template replications from SD 1.4 and interpolated reconstructions (see Section 5.1).
    While the training data for Midjourney remains undisclosed, given the scale of the model and industry norms, we assume that large portions of the training corpus were sourced from publicly available internet data.
    DeepFloyd
    Source Generated Source Generated Source Generated Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    (a) Example results from our attack on DeepFloyd. For each pair, the source image (left) and the generated image (right) are shown.
    Midjourney v4
    Source Generated Source Generated Source Generated Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
    (b) Example results from our attack on Midjourney v4. For each pair, the source image (left) and the generated image (right) are shown.
    Figure 5: Reconstructed images from models from the same generation as SD 1.4
  • SOTA Models: In addition to attacking models from the same generation of SD, we further evaluated the vulnerability of state-of-the-art (SOTA) text-to-image models using collocations identified as template-memorized in SD 1.4. We attacked the following models: Stable Diffusion 3.5 Medium  [5], Flux-Schnell v1.0 [2], and Midjourney v6.1  [3].
    For Stable diffusion and flux Schnell we used the Hugging Face checkpoint in the same manner we’ve conducted our main experiments on SD v1.4, while for Midjourney we used the Discord interface as was detailed previously. We successfully demonstrated some memorization, as seen in Fig. 6, though to a lesser extent than in older models and versions.
    SD 3.5 introduced partial mitigation techniques targeting image-text coupling vulnerabilities; however, these measures were not specifically designed to address the template-style coupling exploited by our attack, which relies on concise prompts derived from external sources rather than training data. Our results indicate that SD 3.5 exhibits increased resilience but remains partially vulnerable to template memorization. Representative examples are provided in Figure 6
    MidJourney V6.1, as discussed, poses additional challenges. However, we applied the same methodology used for V4. While the generated images were typically perturbed (Section 5.2), targeting previously vulnerable benign prompts still yielded images with notable resemblance to the source. For instance, as shown in Fig. 6, the prompt “Abstract Art Essential T-Shirt" (which previously produced near-verbatim copies) continues to generate closely resembling images.
Stable Diffusion v3.5 Flux-Schnell v1.0 MidJourney v6.1
Source Generated Source Generated Source Generated
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Template-memorized images extracted from state-of-the-art models. For each model, the source image (left) and corresponding memorized generation (right) are shown.
4.3 Comparison with previous attacks
As discussed, to allow comparison with previous attacks we also extracted a list of collocations, similar to the one we extracted from the generic websites we chose, by harvesting e-commerce sites that appeared in previous works. For example, we could extract the collocation car seat cover and compare our approach to previous approach that mines captions from the training set. Fig. 7 provide examples of images and the associated prompts used in each attack.
Webster Ours 1 10 1 10 Refer to caption Refer to caption Refer to caption Refer to caption Tribal Aztec Indians pattern Universal Fit Car Seat Covers Floral Car Seat Covers Refer to caption Refer to caption Refer to caption Refer to caption Skull Floor Pillow Size: 36" Floral Floor Pillow Refer to caption Refer to caption Refer to caption Refer to caption 3D Black & White Skull King Design Luggage Covers 007 Floral Luggage Covers
Figure 7: Comparison of images generated after 1 step and 10 steps across prompts from Webster’s attack and ours. Images from top to bottom: (a) Identified by Webster’s white-box attack but not the black-box attack. (b) Identified by the black-box attack as non-verbatim. (c) Identified as template verbatim.
Generated Image Source Image Prompt and Source Caption
Refer to caption Refer to caption prompt: Galaxy Print Universal Fit Car Seat Covers.
Caption: Wild Hearts Can’T Be Broken Car Seat Covers For Horse Lovers 170804 - YourCarButBetter
Refer to caption Refer to caption Prompt: Abstract Art Round Metal Wall Art.
Caption: Designart ’Wide Pathway in Yellow Fall Forest’ Landscape Photo Round Metal Wall Art .
Figure 8: Comparison of Generated and Source Images with Corresponding Prompts and Captions when taking the categories from previous works, examples from [21].
5 Observed Phenomena in Template-Memorized Images
By reconstructing many template-memorized images, we were able to observe the following phenomena, which, to the best of our knowledge, were not previously discussed in literature.
5.1 Interpolation
An important type of image recreation we observed involves interpolation of several real-world images. Several images that were generated by our process were not necessarily verbatim copy of an existing image, nor did they involve naive copying of the background or foreground. Instead, we could find that certain elements were copied from different sources in a single image. This phenomena is demonstrated in Fig. 9. Such type of copying can be challenging to detect. For example, one of the images that can be seen in Fig. 9 copies a tattoo that appeared in a real-world image, but not the whole image. That means that the copied image and the copying image might not be close at all at any metric, and yet it is clear that certain elements in the picture were blatantly copied. This also raises important questions about the level of granularity at which copying should be defined and evaluated. We speculate that some generated images may be interpolated from numerous source images, making them difficult, or even infeasible, to trace with current methods and resources.
Source Images Refer to caption Refer to caption Refer to caption Refer to caption
Generated templates Refer to caption Refer to caption Refer to caption Refer to caption
Interpolations Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Examples of interpolations observed, Left: "X Essential T-Shirt", right: "X iPhone Case & Cover" The elements in the right columns were also identified in the attack of Somepalli et al. [32]
5.2 Perturbations
We’ve noticed clusters of images that were nearly identical in the template sense but with one or more objects perturbed between semantically similar objects in a fixed location such as lamps and chairs. This is exemplified in Fig. 10. Such perturbations preserve a high CLIP similarity (outside the editable region) between the images. Perturbations somewhat affect the pixel-wise $\ell_{2}$ similarity since they change the pixel values, but still maintain a higher-than-random similarity. While difficult to pinpoint the source of this phenomena, it stands in concurrence with the expected consequences of minor perturbations in the noise latent space, as demonstrated in [23]. It might also stem from the one-step denoising behavior described by [36], as a clear layout of the image after the first step, leaves room only for minor changes in the next steps.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Four images with perturbations (see plant on the right) generated from the prompt: “X Wall Tapestry”: where X is: "I Heart ML", "Floral". Rightmost is source found via Google Lens.
5.3 Leakage
Another interesting phenomenon that we identified is a certain template leakage where an image template belonging in one template set appears is generated by a prompt associated with another template group, or not associated with a template at all. In the image generated from template leakage, the object is roughly overlapping with the edited region, in a visually sensible way. Such an example of suspected leakage is demonstrated in Fig. 11(b) where a template associated with the category "T-Shirt" appeared under "Tank Top". The edited area is a tank top, but the background appeared in the source and other generated images with a T-Shirt.
Inspecting e-commerce websites, though, some image templates are re-used for different product. Thus, some cases of suspected leakage may not be such. For example see the common background between "Canvas Wall Art Print" and "Wall Tapestry" Fig. 11(a). The image template appeared with both categories in the training data.
(a) Background shared between template groups, not suspected for leakage. Left: "Galaxy Wall Tapestry", right: "Galaxy Canvas Wall Art Print".
(b) Suspected leakage from "T-Shirt" to "Tank Top". Left: "Abstract Art Essential T-Shirt", right: "Abstract Art Tank Top".
Figure 11: Examples of templates shared between image categories in the training data, including suspected leakage.
Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement FoG-101116258). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work received additional support from the Tel Aviv University Center for AI and Data Science (TAD) and a grant from the Israeli Council of Higher Education.
Appendix A Synthetic Results
We wanted to verify our hypothesis on how the coupling between text templates and image templates causes partial memorization, therefore we conducted synthetic experiments mimicking the sort of coupling in order to intentionally create memorization. Another phenomenon that we wanted to identify and understand is what we term leakage. We have observed on real-world data that certain images from one category included templates from another category. Unfortunately, we could not verify if this is due to leakage from categories or whether certain templates simply appeared on the training data on different category (a phenomenon that we could identify that sometimes happen). Therefore, through synthetic experiments we validate that indeed leakage may be happening.
For our synthetic experiments, we collected 3 images of a coffee Mug with an iPhone SE where we placed the same coffee mug in 3 different locations in our lab. We then manually created a mask for each image with Photopea and then simply replaced the mask area with a pattern using OpenCV. In this stage, the patterns were crudely overlayed, creating an unnatural appearance that was easy for the model to memorize. We observed template verbatim extraction along with interpolation, perturbations and leakage.
As a second stage, we used mockups with more realistic rendering of the overlaid pattern based on the technique used for actual Print on Demand websites. The mockups were taken from Freepik, to one of the mockups we added in Photoshop 2 elements also taken from Freepik: a pair of leaves and a slice of lemon. We selected 3 mockups and replaced the smart object contents in Photoshop by an automation script with some manual fixes to ensure that the overlay is uniform between the images. Beside the different templates, the rest of the experiment remained the same as the first stage. We did not observed template verbatim extraction, but we did observe interpolation, perturbations, and object memorization.
To validate the coupling between text and images, we divide this experiment into two stages. In each stage of this experiment we created a dataset mimicking the data from design placement websites on images created by us for that purpose such that they didn’t appear in the training set. We selected the category of "Coffee Mug" which did not involve identified templates. Next, we select 3 images of a blank coffee mug and overlay patterns on them as detailed below. The overlaid patterns are 18 patterns with CC0 license downloaded from Freepik [7]. The caption for each image is composed of a short description for the pattern e.g. "Floral" and the category "Coffee Mug". The dataset of 54 image-caption pairs were used to fine-tune Stable Diffusion by LoRa [22]. We then generated images from the trained model using the prompt template “skg Coffee Mug” where “skg” had been replaced with a pattern description, including patterns that were not in the training set.
In a addition to generating the collocation "Coffee Mug", we generate images from our fine tuned model through semantically similar and semantically dissimilar to “Coffee Mug”: “Tea Cup” and “T-Shirt”, Accordingly. Through these additional prompts we could identify that leakage indeed happens. Specifically for the prompts "skg Tea Cup", the backgrounds of all of the images are templates interpolation or at least similar colors/textures, the tea cups are also in top view which was not seen during training and did not appear in the coffee mug generated images (only the 3 views from the train set), and also in different shapes.
The results of our synthetic experiments support our assumption on how the template text-image structure in the training set encourages template-memorization, as well as demonstrating the observed phenomena we discussed at 5.
stage I Stage II
Training data Refer to caption Refer to caption
"Coffee Mug" Refer to caption Refer to caption
"Tea Cup" Refer to caption Refer to caption
"T-Shirt" Refer to caption Refer to caption
Figure 12: Intentionally causing template memorization by fine tuning SD on coupled image-text pairs. The generated results demonstrate the phenomena of interpolation, perturbations, and leakage.
Appendix B Traces of memorization in Stable Diffusion 3.5
Beside clear template extraction up to perturbations, we also noticed that even in cases where our original attack failed, traces of the attacks could still be found and perhaps utilized in future, more refined, attacks. For example, although the beach towel in Fig. 3 could no longer be reconstructed directly, identifying its source via Google Lens revealed that the caption still produces distorted versions of the original image. This indicates that training images from e-commerce websites may continue to pose risks in future iterations of diffusion models.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Traces of our attack in SD 3.5 (Medium). The first three images were generated using the prompt "BOYOUTH Round Beach Towel,X Beach Mat with Tassels Ultra Soft Super Water Absorbent Multi-Purpose Towel,59 inch-Diameter" Taken from the source image’s product description in Amazon, where X was: "Abstract Art", "Floral", "Galaxy". The image to the right of the vertical line is the source image, which was reconstructed in previous models (see Fig. 3)